Julie invested 5355 dollars in a stock. the first 7 years it went down by 22 percent each year. The following 9 years it went up by 38 percent each year. How much money does Julie have at the end?
Answer by Theo(13342)

$5355 is the initial investment.
the formula to use is:
f = p * (1+i)^n
f = future value
p = present value
i = annual interest rate
n equal number of years.

for the first 7 years the equation becomes:
p = 5355
i = -.22
n = 7
f = 5355 * (1-.22)^7
use your calculator to get:
f = 940.6362122

for the next 9 years the equation becomes:
p = 940.6362122
i = .38
n = 9
f = 940.6362122 * (1.38)^9
use your calculator to get:
f = 17073.92902